A+b+c+d+e=21
A+b=6
B+c=10
C+d=7
D+e=9

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52786)

.
a + b + c + d + e = 21    (1)

a + b = 6                 (2)

b + c = 10                (3)

c + d = 7                 (4)

d + e = 9                 (5)



1.  From equations (2) and (5), substitute  a + b = 6  and   d + e = 9  into equation (1). You will get

    6 + c + 9 = 21,  which implies  c = 21 - 6 - 9 = 6.



2.  Now equation   (3) becomes  b + 6 = 10,  giving  b = 10 - 6 = 4,

    while equation (4) becomes  6 + d =  7,  giving  d =  7 - 6 = 1.



3.  Now equation   (2) becomes  a + 4 = 6,  giving  a = 6 - 4 = 2,

    while equation (5) becomes  1 + e = 9,  giving  e = 9 - 1 = 8.



Answer.  a= 2,  b= 4,  c= 6,  d= 1  and  e= 8.

Solved.


Answer by greenestamps(13200)



Adding equations (2) and (4) gives you a+b+c+d=13; comparing that to equation (1) gives you e=8. Then that with equation (5) gives you d=1.

Adding equations (2) and (5) gives you a+b+d+e=15; comparing that to equation (1) gives you c=6. Then that with equation (3) gives you b=4; and then that with equation (2) gives you a=2.

ANSWER: a=2; b=4; c=6; d=1; e=8